# ERASE: Benchmarking Feature Selection Methods for Deep Recommender Systems

### v2

@@ Abstract.
Deep Recommender Systems (DRS) are increasingly dependent on a large number of feature fields for more precise recommendations. Effective feature selection methods are consequently becoming critical for further enhancing the accuracy and optimizing storage efficiencies to align with the deployment demands. This research area, particularly in the context of DRS, is nascent and faces three core challenges. Firstly, variant experimental setups across research papers often yield unfair comparisons, obscuring practical insights. Secondly, the existing literature’s lack of detailed analysis on selection attributes, based on large-scale datasets and a thorough comparison among selection techniques and DRS backbones, restricts the generalizability of findings and impedes deployment on DRS. Lastly, research often focuses on comparing the peak performance achievable by feature selection methods, an approach that is typically computationally infeasible for identifying the optimal hyperparameters and overlooks evaluating the robustness and stability of these methods. To bridge these gaps, this paper presents ERASE, a comprehensive bEnchmaRk for feAture SElection for DRS. ERASE comprises a thorough evaluation of eleven feature selection methods, covering both traditional and deep learning approaches, across four public datasets, private industrial datasets, and a real-world commercial platform, achieving significant enhancement. Our code is available online [^1] https://github.com/Applied-Machine-Learning-Lab/ERASE for ease of reproduction.
@@
* Authors contributed equally to this research.
copyright: acmcopyrightjournalyear: 2024conference: 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; AUG 25 - 29, 2024; Barcelona, Spain
@@ 1. Introduction
To improve the accuracy of recommendations, DRS is progressively integrating an expanding array of feature fields into their predictive models, which can number in the hundreds or even thousands (Heaton, 2016). The significance of each feature varies, resulting in the accumulation of superfluous or extraneous features. Consequently, feature selection, which concentrates on pinpointing and leveraging the most critical features, is becoming increasingly crucial in modern DRS (Zheng et al., 2023; Chen et al., 2022). One immediate benefit of feature selection is the enhancement of prediction performance, achieved by eliminating non-contributory features that could otherwise adversely affect predictions. From an industrial viewpoint, selecting predictive features is also essential for meeting deployment criteria regarding memory usage since unnecessary storage demands can often be inflated by the presence of redundant features.
@@
Hand-crafted featrue selection usually requires lots of expert knowledge and labor efforts, which is usually infeasible to achieve the optimal results when the candidate set contains thousands of features. Researchers have invised various methods to automatically select predictive features, including statistical methods (Tibshirani, 1996; Liu and Setiono, 1995; Vergara and Estévez, 2014), learning methods (Chen and Guestrin, 2016; Friedman, 2001; Breiman, 2001), and agent-based methods (Liu et al., 2019; Fan et al., 2020; Zhao et al., 2020).
@@
Despite these methods achieving good results in feature selection, they also reveal the following issues that hinder the development of this field. 1) Experimental Differences. Recent years have witnessed a variety of methods designed for DRS (Wang et al., 2022; Lin et al., 2022; Guo et al., 2022; Lee et al., 2023; Lyu et al., 2023; Zhang et al., 2023; Wang et al., 2023). However, these works conduct experiments with varied settings. For example, MultiSFS (Wang et al., 2023) is designed to select features for multi-taske DRS. SHARK (Zhang et al., 2023) suggests the feature selection method F-permutation together with a quantization method. Optfs (Lyu et al., 2023) selects the feature from the value-level while others mainly select from the field-level. These experimental differences usually lead to the unfair or unavailable comparison, making the subsequent researchers stuggling to generate practical insights. 2) Insufficiency. Existing feature selection benchmarks are predominantly tailored for conventional downstream tasks, such as classification (Bolón-Canedo et al., 2014; Bommert et al., 2020). They are primarily built upon synthetic or domain-specific datasets (Bolón-Canedo et al., 2013; Forman, 2003; Darshan and Jaidhar, 2018), which diverges significantly from the complex, large-scale datasets encountered in DRS. This divergence results in a notable deficiency in guiding feature selection specifically for DRS, due to the benchmarks’ disparate scope and focus. DeepLasso (Cherepanova et al., 2023), while being among the benchmarks most closely related to DRS, primarily addresses tabular learning—a context that, despite similarities, does not fully align with the intricacies of recommendation tasks. Furthermore, such benchmarks often limit their exploration to traditional selection methods and rely on datasets of a much smaller scale than those typically associated with DRS, thereby omitting crucial, in-depth analysis from an industrial perspective. Moreover, while related literature surveys (Chen et al., 2022; Zheng et al., 2023) provide comprehensive reviews of the field, they fall short in offering empirical evidence or experimental results that could motivate practical deployments in DRS, lacking the necessary data-driven support to inform and inspire future research directions. 3) Assessment Deficit. A pivotal hyperparameter in the realm of feature selection is the number of features to be selected, denoted as $k$ in this study. Feature selection methods often exhibit variability in performance across different values of $k$, and it is common that the optimal $k$ varies from one method to another (Wang et al., 2022; Lin et al., 2022; Lyu et al., 2023). This variability introduces two significant challenges in evaluating feature selection methods comprehensively. First, the direct comparison of methods at their respective optimal $k$ may not constitute a fair assessment. Such comparisons fail to account for the different memory requirements associated with varying optimal $k$ values for different selection methods. Furthermore, this approach neglects the performance variability under sub-optimal hyperparameter settings, thereby obscuring insights into the methods’ robustness and stability across a range of $k$ values. Second, the exhaustive search for the optimal $k$ across the entire spectrum of possible values is both time-consuming and computationally intensive. This necessitates an evaluation methodology capable of effectively assessing a method’s performance based on partial $k$ values, offering a more efficient means to gauge feature selection effectiveness without finishing complete iterations.
@@
  • We present ERASE, a comprehensive benchmark for DRS feature selection methods, providing the fair comparison for emerging selection techniques with various datasets and DRS backbones. To the best of our knowledge, we are the first to focus on benchmarking feature selection methods for recoomendation tasks.
@@ Methods
In this section, we will give an overview of our benchmark. As shown in Figure 1, the benchmark is consisted of four components: dataset, feature selection methods, backbone models, and metrics. Specifically, datasets are crucial experimental foundations for benchmarks. We elaborate on the datasets utilized in Section 2.1. The core introduction to the feature selection methods is presented in Section 2.2. Next, different feature selection methods or the selected features need to apply to specific backbone models, which are discussed in Section 2.3. Finally, in Section 2.4, we provide a detailed description of the evaluation metrics used in this benchmark, including two existing metrics and one newly proposed metric.
@@ 2.1. Datasets
To comprehensively evaluate the effectiveness of different feature selection methods, we select four public datasets for our experiments. Avazu [^2] https://www.kaggle.com/competitions/avazu-ctr-prediction and Criteo [^3] https://ailab.criteo.com/ressources/ are selected because they are frequently used dataset for studying feature selection in DRS (Wang et al., 2022; Lin et al., 2022; Lyu et al., 2023). To compare the performance of different methods with a small feature set, we choose the popular Movielens-1M [^4] https://grouplens.org/datasets/movielens/1m/ dataset in the recommendation field. Additionally, to compare the performance of different methods in scenarios closer to real-world recommender systems, we supplement with the AliCCP [^5] https://tianchi.aliyun.com/dataset/408 dataset, which possesses user and item features and includes a rich set of 85,316,519 interaction samples. The statistics of datasets and the detailed introduction of datasets are illustrated in Appendix A.
@@ 2.2. Feature Selection Methods
Figure 2. Overview of three categories of feature selection methods in deep recommender systems.
@@
We list all feature selection methods in this work and their attributes in Table 1. There are three feasible dimensions to classify these methods. 1) Training strategy. Based on the training strategy, the methods can be single-stage or two-stage. Single-stage usually integrates the feature selection module directly into the original model without changing the training logic. In contrast, two-stage methods contain searching and retraining phases. Informative feature fields or feature values are selected in the searching phase, and the backbone model is trained with these selected features in the retraining phase. 2) Selection type. There are two types of selection in feature selection methods: soft selection and hard selection. The soft selection offers a mask to affect inputs. Features multiplied with 0 are considered filtered out. For the hard selection, features are removed directly from the inputs. 3) Selection technique. Depending on the techniques used for selection, as shown in Figure 2, we divide methods contained in our benchmark into three categories: shallow feature selection, gate-based selection, and sensitivity-based selection. Shallow selection methods, including Lasso (Tibshirani, 1996), GBDT (Friedman, 2001), RandomForest (RF) (Breiman, 2001) and XGBoost (Chen and Guestrin, 2016), select features based on the feature importance score derived from the shallow learning methods (e.g., RandomForest). Gate-based methods, including AutoField (Wang et al., 2022), AdaFS (Lin et al., 2022), OptFS (Lyu et al., 2023), and LPFS (Guo et al., 2022), distribute gates to control the usage of features and decide the selection results according to gate outputs. Sensitivity-based methods, including Permutation (Fisher et al., 2019), SHARK (Zhang et al., 2023), and SFS (Wang et al., 2023), compute the feature importance by measuring the sensitivity of gradients to features. Due to the unbalanced distribution of feature selection methods classifying on training strategy and selection type, we elaborate our work based on selection technique classification. The detailed introduction of these methods is in Appendix B.
@@ 2.3. Backbone Models
To comprehensively evaluate the effectiveness of feature selection methods, we choose four popular DRS models as backbones: 1) Wide&Deep: A classical model contains shallow and deep networks to capture feature interactions. 2) DeepFM: A model with FM module to automatically learn feature interactions. 3) DCN: A model with cross layers to study feature interactions. 4) FibiNet: A model equipped with SENet and bilinear layer to adaptively learns feature importance and captures high-order feature interactions. The detailed introduction of these models are in Appendix C.
@@ 2.4. Metrics
In this paper, we focus on the Click-Through Rate (CTR) prediction task, so we take AUC and Logloss as the two main metrics in our benchmark, the detailed introduction is in Appendix E. Additionally, since there is currently no metric that comprehensively evaluates the performance of feature selection methods across varying numbers of selected features, we propose a new evaluation metric named “AUKC” to address this challenge.
@@
where $|K|$ denotes the number of all feature fields, $AUC_{k}$ is the AUC score when selecting $k$ feature fields follow a specific feature selecting method. If there is no input features, the model would make random predictions so that $AUC_{0}=0.5$.
In practice, due to resource contrains, it is generally not feasible to conduct experiments with every possible number of selections. Therefore, we further propose a more general form of AUKC to accommodate this change:
@@
where $|K|$ denotes the number of all feature fields, $|N|$ is the number of segments across the entire length of the feature fields set. $AUC_{n,l}$ is the AUC corresponding to the number of features selected at the left endpoint of the $n$-th segment, and $AUC_{n,r}$ denotes the AUC with the number of selections at the right endpoint of the $n$th segment. $\Delta l_{n}$ represents the length of the $n$-th segment. The detailed process of the formula derivation is in the Appendix F.
AUKC and AUC share a similar conceptual basis, representing the area under a curve, and both have a value range from 0 to 1. However, unlike AUC, the endpoint of AUKC does not necessarily equal 1, and the values in the middle of the curve can exceed the endpoint value. This occurs because eliminating features with less information can result in a model that performs better than one using all available features. The AUKC metric takes into account the effectiveness of feature selection at different numbers of selected features, $k$, thereby providing a more comprehensive reflection of the efficacy of feature selection methods.
@@ 3. Experiments
In this section, the following research questions will be answered:
@@
RQ2: How do different feature selection methods perform in terms of robustness and stability across various numbers of selections?
RQ3: How is the efficiency of different feature selection methods?
@@
RQ5: Do the performances of various feature selection methods remain consistent online and offline?
Importantly, we also elucidate the congruence among feature rankings derived from various selection methods by visualizing their similarities in Appendix J. This analysis aims to explore both the intra- and inter-group relations among the three categories.
@@ 3.1. Experimental Details
We implement the benchmark framework based on Pytorch 1.11. In the training phase, we utilize the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-8}$. We set the learning rate as 0.001, the batch size as 4,096, and the embedding size as 8. For the activation function, if the original papers do not emphasize a specific one, we use ReLU as the activation function. We release the repository of our benchmark online [^6] https://anonymous.4open.science/r/ERASE-B28E. The details of reproduction and hyperparameters for specific backbone models and feature selection methods are illustrated in Appendix D and Appendix G. We conduct each experimental setting three times and record its average metrics to mitigate the impact of experimental fluctuations.

### v3

@@ Abstract.
Deep Recommender Systems (DRS) are increasingly dependent on a large number of feature fields for more precise recommendations. Effective feature selection methods are consequently becoming critical for further enhancing the accuracy and optimizing storage efficiencies to align with the deployment demands. This research area, particularly in the context of DRS, is nascent and faces three core challenges. Firstly, variant experimental setups across research papers often yield unfair comparisons, obscuring practical insights. Secondly, the existing literature’s lack of detailed analysis on selection attributes, based on large-scale datasets and a thorough comparison among selection techniques and DRS backbones, restricts the generalizability of findings and impedes deployment on DRS. Lastly, research often focuses on comparing the peak performance achievable by feature selection methods. This approach is typically computationally infeasible for identifying the optimal hyperparameters and overlooks evaluating the robustness and stability of these methods. To bridge these gaps, this paper presents ERASE, a comprehensive bEnchmaRk for feAture SElection for DRS. ERASE comprises a thorough evaluation of eleven feature selection methods, covering both traditional and deep learning approaches, across four public datasets, private industrial datasets, and a real-world commercial platform, achieving significant enhancement. Our code is available online [^1] https://github.com/Applied-Machine-Learning-Lab/ERASE for ease of reproduction.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671571isbn: 979-8-4007-0490-1/24/08ccs: Information systems Recommender systems
@@ 1. Introduction
To improve the accuracy of recommendations, DRS is progressively integrating an expanding array of feature fields into their predictive models, which can number in the hundreds or even thousands (Heaton, 2016). The significance of each feature varies, resulting in the accumulation of superfluous or extraneous features. Consequently, feature selection, which concentrates on pinpointing and leveraging the most critical features, is becoming increasingly crucial in modern DRS (Zheng et al., 2023; Chen et al., 2024). One immediate benefit of feature selection is the enhancement of prediction performance, achieved by eliminating non-contributory features that could otherwise adversely affect predictions. From an industrial viewpoint, selecting predictive features is also essential for meeting deployment criteria regarding memory usage since unnecessary storage demands can often be inflated by the presence of redundant features.
@@
Hand-crafted feature selection usually requires lots of expert knowledge and labor efforts, which is usually infeasible to achieve optimal results when the candidate set contains thousands of features. Researchers have designed various methods to automatically select predictive features, including statistical methods (Tibshirani, 1996; Liu and Setiono, 1995; Vergara and Estévez, 2014), learning methods (Chen and Guestrin, 2016; Friedman, 2001; Breiman, 2001), and agent-based methods (Liu et al., 2019; Fan et al., 2020; Zhao et al., 2020). Despite the satisfactory results these methods achieved, they also reveal the following issues hindering the development of this field:
  • Experimental Differences. Recent years have witnessed a variety of methods designed for DRS (Wang et al., 2022; Lin et al., 2022; Guo et al., 2022; Lee et al., 2023; Lyu et al., 2023; Zhang et al., 2023; Wang et al., 2023). However, these works conduct experiments with varied settings. For example, MultiSFS (Wang et al., 2023) is designed to select features for multi-task DRS. SHARK (Zhang et al., 2023) suggests the feature selection method F-permutation and a quantization method. Optfs (Lyu et al., 2023) selects the feature from the value-level while others mainly select from the field-level. These experimental differences usually lead to unfair or unavailable comparisons, making the subsequent researchers struggle to generate practical insights.
@@
  • Insufficiency. Existing feature selection benchmarks are predominantly tailored for conventional downstream tasks, such as classification (Bolón-Canedo et al., 2014; Bommert et al., 2020). They are primarily built upon synthetic or domain-specific datasets (Bolón-Canedo et al., 2013; Forman, 2003; Darshan and Jaidhar, 2018), which diverges significantly from the complex, large-scale datasets encountered in DRS. This divergence results in a notable deficiency in guiding feature selection specifically for DRS, due to the benchmarks’ disparate scope and focus. DeepLasso (Cherepanova et al., 2023), while being among the benchmarks most closely related to DRS, primarily addresses tabular learning—a context that, despite similarities, does not fully align with the intricacies of recommendation tasks. Furthermore, such benchmarks often limit their exploration to traditional selection methods and rely on datasets of a much smaller scale, thereby omitting crucial, in-depth analysis from an industrial perspective for DRS. Moreover, while related literature surveys (Chen et al., 2024; Zheng et al., 2023) provide comprehensive reviews of the field, they fall short in offering empirical evidence or experimental results that could motivate practical deployments in DRS, lacking the necessary data-driven support to inform and inspire future research directions.
  • Assessment Deficit. A pivotal hyperparameter for feature selection methods is the number of features to be selected, denoted as $k$ in this study. Feature selection methods often exhibit significant variability in performance across different values of $k$, and the optimal $k$ is not the same for all methods (Wang et al., 2022; Lin et al., 2022; Lyu et al., 2023). This variability on $k$ introduces two significant challenges in evaluating feature selection methods. First, the direct comparison of methods at their respective optimal $k$ may not constitute a fair assessment. Such comparisons fail to account for the different memory requirements associated with varying optimal $k$ values for different selection methods. Furthermore, this approach neglects the performance variability under sub-optimal hyperparameter settings, thereby obscuring insights into the methods’ robustness and stability across a range of $k$ values. Second, the exhaustive search for the optimal $k$ across the entire spectrum of possible values is time-consuming and computationally intensive. This necessitates an evaluation methodology capable of effectively assessing a method’s performance based on partial $k$ values, offering a more efficient means to gauge feature selection effectiveness without finishing complete iterations.
@@
  • We present ERASE, a comprehensive benchmark for DRS feature selection methods, providing a fair comparison for emerging selection techniques with various datasets and DRS backbones. To the best of our knowledge, we are the first to focus on benchmarking feature selection methods for recommendation tasks.
@@ Methods
In this section, we will give an overview of our benchmark. As shown in Figure 1, the benchmark consists of four components: dataset, feature selection methods, backbone models, and metrics.
@@ 2.1. Datasets
To comprehensively evaluate the effectiveness of different feature selection methods, we select four public datasets for our experiments. Avazu [^2] https://www.kaggle.com/competitions/avazu-ctr-prediction and Criteo [^3] https://ailab.criteo.com/ressources/ are selected because they are frequently used dataset for studying feature selection in DRS (Wang et al., 2022; Lin et al., 2022; Lyu et al., 2023). To compare the performance of different methods with a small feature set, we choose the popular Movielens-1M [^4] https://grouplens.org/datasets/movielens/1m/ dataset in the recommendation field. Additionally, to compare the performance of different methods in scenarios closer to real-world recommender systems, we supplement with the AliCCP [^5] https://tianchi.aliyun.com/dataset/408 dataset, which possesses user and item features and includes a rich set of 85,316,519 interaction samples. The statistics of datasets and the detailed introduction are illustrated in Appendix A.
@@ 2.2. Feature Selection Methods
Figure 2. Overview of three categories of feature selection methods in deep recommender systems. Shallow methods typically use statistical algorithms to assign feature importance to each field. Gate-based methods, on the other hand, assign gates to feature embeddings and consider the gate values as the importance of the corresponding features. Sensitivity-based methods derive parameter sensitivity from backward propagation steps and calculate feature importance accordingly.
We list all feature selection methods in this work and their attributes in Table 1. There are three feasible dimensions to classify these methods. 1) Training strategy. Based on the training strategy, the methods can be single-stage or two-stage. Single-stage usually directly integrates the feature selection module into the original model without changing the training logic. In contrast, two-stage methods contain searching and retraining phases. Informative feature fields or feature values are selected in the searching phase, and the backbone model is trained with these selected features in the retraining phase. 2) Selection type. There are two types of selection in feature selection methods: soft selection and hard selection. The soft selection offers a mask to affect inputs. Features multiplied with 0 are considered filtered out. For the hard selection, features are removed directly from the inputs. 3) Selection technique. Depending on the techniques used for selection, as shown in Figure 2, we divide methods contained in our benchmark into three categories: shallow feature selection, gate-based selection, and sensitivity-based selection.
Due to the unbalanced distribution of feature selection methods classifying on training strategy and selection type, we elaborate our work based on selection technique classification. Specifically, our benchmark contains the following methods:
1. Shallow Feature Selection
  • GBDT (Friedman, 2001). Gradient-boosted decision tree (GBDT) achieves superior performance by continually adding trees to fit residuals. By aggregating the feature importance scores from each tree, it also serves as an effective method for feature selection.
  • RandomForest (Breiman, 2001) (RF). RandomForest derives the feature importance by measuring how much each feature decreases the impurity in a tree, commonly using Gini impurity or entropy.
  • XGBoost (Chen and Guestrin, 2016). Extreme Gradient Boosting (XGBoost) ranks the importance of features by calculating the improvement of every feature on the final performance.
@@ 1. Shallow Feature Selection
2. Gate-based Feature Selection
  • AutoField (Wang et al., 2022). AutoField is the first gate-based feature selection method in DRS. It designs a novel controller network to generate a 2-dimensional vector determining whether to choose this feature field. It is a two-stage method and selects features on the field level.
  • LPFS (Guo et al., 2022). LPFS argues that the conclusion that features with smaller weights are less important than those with larger weights may not be correct. It proposes a kind of smoothed-$l^{0}$ function that can effectively select informative features. It is a single-stage method and focuses on the field level.
3. Sensitivity-based Feature Selection
@@ 2.3. Backbone Models
To comprehensively evaluate the effectiveness of feature selection methods, we choose four popular DRS models as backbones: 1) Wide&Deep: A classical model contains shallow and deep networks to capture feature interactions. 2) DeepFM: A model with FM module to automatically learn feature interactions. 3) DCN: A model with cross layers to study feature interactions. 4) FibiNet: A model equipped with SENet and bilinear layer to adaptively learn feature importance and capture high-order feature interactions. The detailed introduction of these models is in Appendix B.
@@ 2.4. Metrics
In this paper, we focus on the Click-Through Rate (CTR) prediction task, so we take AUC and Logloss as the two main metrics in our benchmark. The detailed introduction is in Appendix C. Additionally, since there is currently no metric that comprehensively evaluates the performance of feature selection methods across varying numbers of selected features, we propose a new evaluation metric named “AUKC” to address this challenge.
@@
where $|K|$ denotes the number of all feature fields, $AUC_{k}$ is the AUC score when selecting $k$ feature fields follow a specific feature selecting method. If there are no input features, the model would make random predictions so that $AUC_{0}=0.5$.
In practice, due to resource constraints, it is generally not feasible to conduct experiments with every possible number of selections. Therefore, we further propose a more general form of AUKC to accommodate this change:
@@
where $|K|$ denotes the number of all feature fields and $|N|$ is the number of segments across the entire length of the feature fields set. $AUC_{n,l}$ is the AUC corresponding to the number of features selected at the left endpoint of the $n$-th segment, and $AUC_{n,r}$ denotes the AUC with the number of selections at the right endpoint of the $n$th segment. $\Delta l_{n}$ represents the length of the $n$-th segment. The detailed process of the formula derivation is in Appendix D.
AUKC and AUC share a similar conceptual basis, representing the area under a curve, and both have a value range from 0 to 1. However, unlike AUC, the endpoint of AUKC does not necessarily equal 1, and the values in the middle of the curve can exceed the endpoint value. This occurs because eliminating features with less information can result in a model that performs better than one using all available features. The AUKC metric considers the effectiveness of feature selection at different numbers of selected features, $k$, thereby providing a more comprehensive reflection of the efficacy of feature selection methods.
@@ 3. Experiments
In this section, we will provide extensive experimental results to answer following research questions:
@@
  • RQ2: How do feature selection methods perform in terms of robustness and stability across various numbers of selections?
  • RQ3: How is the efficiency of feature selection methods?
@@
  • RQ5: Do the performances of feature selection methods remain consistent online and offline?
  • RQ6: What is the relationship between feature rankings of different feature selection methods?
@@ 3.1. Experimental Details
We implement the benchmark framework based on Pytorch 1.11. In the training phase, we utilize the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-8}$. We set the learning rate as 0.001, the batch size as 4,096, and the embedding size as 8. For the activation function, if the original papers do not emphasize a specific one, we use ReLU as the activation function. We release the repository of our benchmark online [^6] https://github.com/Applied-Machine-Learning-Lab/ERASE. We conduct each experimental setting three times and record its average metrics to mitigate the impact of experimental fluctuations.